OpenStudy (wach):

2x^4y^-3x^-19 + y^1/3y^-3/4 Can someone check this for me? I'm not confident in my answer. The answer I got is 2/y^3x^15 + the square root of y.

OpenStudy (anonymous):

everything looks correct, as long as its \[\frac{ 2 }{ x ^{15}y ^{3} } + \frac{ 1 }{ \sqrt{y} }\]

OpenStudy (wach):

Oh, yes ^ Thank you! :)

OpenStudy (anonymous):

np :)